OpenStudy (anonymous):

Solve this system of equations using any method. 2x – y = 5 4x + y = 7 (4, –3) (4, –9) (2, –1) (2, –4) :)

OpenStudy (anonymous):

Add the two functions together to determine x.

OpenStudy (anonymous):

http://www.mathpapa.com/algebra-calculator.html here it shows you how

OpenStudy (alexandervonhumboldt2):

y=2y-5 y=7-4x

OpenStudy (alexandervonhumboldt2):

7-4x=2x-5

OpenStudy (alexandervonhumboldt2):

12=6x x=2

OpenStudy (alexandervonhumboldt2):

y=-1

OpenStudy (alexandervonhumboldt2):

(2, -1)

OpenStudy (anonymous):

That is what i thought. Thank you!

OpenStudy (anonymous):

Each small kite costs $4, while each large kite costs $7. If the total money collected selling kites is $150, and 30 kites are sold, which system of equations can be used to find out how many of each size kite are sold? Let s be the small kites and l be the number of large kites.

OpenStudy (anonymous):

What about that one?

OpenStudy (alexandervonhumboldt2):

4s+7l=150 l+s=30

OpenStudy (alexandervonhumboldt2):

l=30-s 4s+7*(30-s)=150